OpenStudy (anonymous):

Add.7 3/10+6 1/3+2 7/10

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

Where are you struggling?

OpenStudy (triciaal):

are you just adding mixed numbers? 7+3/10 +6 +1/3 +2 +7/10 add the whole numbers 7 + 6 + 2 = add the fractions 3/10 + 7/10 =10/10 =1 add the whole numbers and fractions what is your answer?

OpenStudy (anonymous):

i got it